### v1

# Bayesian Rank-Clustering

Abstract
In a traditional analysis of ordinal comparison data, the goal is to infer an overall ranking of objects from best to worst with each object having a unique rank. However, the ranks of some objects may not be statistically distinguishable. This could happen due to insufficient data or to the true underlying abilities or qualities being equal for some objects. In such cases, practitioners may prefer an overall ranking where groups of objects are allowed to have equal ranks or to be rank-clustered. Existing models related to rank-clustering are limited by their inability to handle a variety of ordinal data types, to quantify uncertainty, or by the need to pre-specify the number and size of potential rank-clusters. We solve these limitations through the proposed Bayesian Rank-Clustered Bradley-Terry-Luce model. We allow for rank-clustering via parameter fusion by imposing a novel spike-and-slab prior on object-specific worth parameters in Bradley-Terry-Luce family of distributions for ordinal comparisons. We demonstrate the model on simulated and real datasets in survey analysis, elections, and sports.
Keywords ordinal comparisons  $\cdot$ Plackett-Luce  $\cdot$ Bradley-Terry  $\cdot$ spike-and-slab  $\cdot$ fusion priors  $\cdot$ item indifference
1 Introduction
In a traditional analysis of ordinal data, we assume $I$ judges assess $J$ objects by providing ordinal preferences, $\Pi$. Each judge’s ordinal preferences, $\Pi_{i}$, may be provided in various forms, such as complete rankings, partial rankings, or pairwise comparisons among the available objects or some subset thereof. Standard statistical model families for ranking data (e.g., Mallows (Mallows, 1957), Bradley-Terry-Luce (Bradley and Terry, 1952; Plackett, 1975; Luce, 1959)) derive or estimate the rank of each object such that each object receives a unique rank. An overall ranking orders all objects from best to worst on the basis of their estimated rank. Analyses of this kind are used to rank candidates in an election given ranked choice votes (Gormley and Murphy, 2008; Mollica and Tardella, 2017), sports teams or players in a league using game outcomes (Tutz and Schauberger, 2015; Barrientos et al., 2023), or genes based on ordinal comparisons of genomic data (Eliseussen et al., 2023; Vitelli et al., 2018).
However, requiring estimated ranks to be unique is not always useful or appropriate. Specifically, some objects may be equal or indistinguishable in quality or ability at the population level. In such cases, we refer to the objects as being rank-clustered. For example, consider one situation of an election in which 2 candidates, both of the same political party, are running for an office. If voters express their preferences solely on the basis of party, the candidates are inherently equal in quality; the rational voter would randomly select a candidate to vote for. In another case, when the number of votes cast is small, estimated ranks assigned to each candidate could exhibit substantial uncertainty, suggesting the candidates are indistinguishable in quality based on the limited number of observed votes. In such situations, allowing for rank-clusters may improve interpretability and aid accurate inference, prediction, and decision-making when analyzing ordinal preferences.
In this paper, we propose a Bayesian framework for ordinal data analysis that estimates an overall ranking of objects with rank-clusters. Specifically, we choose to model observed rank via the Bradley-Terry-Luce (BTL) family of distributions which permits analysis of ordinal preferences in many forms, such as complete rankings, partial rankings, pairwise comparisons, and groupwise comparisons. To induce rank clusters, we place a novel spike-and-slab prior that partitions continuous parameters, proposed herein, on the object-specific parameters of BTL distributions. In contrast to existing work related to rank-clustering in the literature, our model does not require the parameter order nor the number or size of rank-clusters to be known in advance. Instead, these quantities are treated as random variables and estimated simultaneously, and their corresponding uncertainty is included in the resulting inferences. In this paper, we develop a computationally-efficient Gibbs sampler for estimation and apply the model to real and simulated data.
The rest of this work is organized as follows. We first review relevant literature related to rank-clustering in Section 2. Then, we propose the Partition-based Spike-and-Slab Fusion prior and apply it to a BTL model for ordinal data in Section 3. We develop a computationally-efficient Gibbs sampler based on reversible jump Markov chain Monte Carlo in Section 4 and demonstrate its accuracy on simulated data. To demonstrate a wide variety of methodological values of our proposed framework, Section 5 applies the model to four real datasets: (i) complete rankings of sushi preferences provided by Japanese adults, (ii) partial rankings of 2021 Minneapolis mayoral candidates expressed by voters in a ranked choice election, (iii) complete and partial rankings of policy options from Eurobarometer 34.1, a survey which measures various European attitudes, and (iv) pairwise basketball game outcomes from the 2020-21 season of the National Basketball Association. We conclude with a brief discussion in Section 6.
2 Background
Before considering existing literature related to this paper, it is helpful to review some basic terminology and notation for ranking data. Rankings are a type of ordinal preference data that denote a relative ordering of objects from best to worst (potentially allowing ties). We use the operator $\prec$ to denote a strict ordering of two objects; e.g., $A\prec B$ states that object $A$ is strictly preferred to $B$. In a ranking, an object’s rank is the place it receives in the ranking. [^1] Although some authors have drawn a distinction between the terms “ranking” and “ordering,” in this paper we choose to use solely the former in accordance with its popular usage. Rankings arise in different forms. Given a collection of objects, a ranking is called complete when all objects are ranked. In contrast, a ranking is called partial when all objects were considered, but only a subset of the most-preferred are ranked (e.g., a top-$5$ ranking). In a partial ranking, we assume that unranked objects are less-preferred than those ranked, but also that the preference order among the unranked objects is unknown. Next, we call a ranking incomplete when a judge is asked only to rank a subset of the complete collection of objects. In incomplete rankings, no information can be gleaned regarding objects not considered. For example, if a voter is asked by an election pollster to rank candidates from a single political party, the ranking should provide no information regarding their preferences on candidates from other parties. We call incomplete rankings involving two objects (candidates in the above example) a pairwise comparison, and incomplete rankings involving more than two objects a groupwise comparison. Rankings may be both partial and incomplete; e.g., a top-3 ranking of mayoral candidates only among the 5 candidates from a specific political party.
In the rest of this section, we will briefly review methods for estimating rank-clusters based on the Bradley-Terry-Luce and Mallows families of ordinal data models in turn. For a more thorough review of these standard model families, see e.g., Marden (1996); Alvo and Yu (2014).
2.1 Methods based on Bradley-Terry-Luce Distributions
Most work related to rank-clustering utilized the Bradley-Terry-Luce (BTL) family, which comprises the Bradley-Terry and Plackett-Luce distributions and their extensions. The Bradley-Terry model, proposed by Zermelo (1929) and discovered independently by Bradley and Terry (1952), is parameterized by the vector $\omega\in\mathbb{R}_{>0}^{J}$, in which each $\omega_{j}$ corresponds to the worth of object $j$. Specifically, the Bradley-Terry model specifies the probability that object $i$ will be ranked above object $j$ in pairwise tournament as
$P[i\prec j|\omega_{i},\omega_{j}]=\frac{\omega_{i}}{\omega_{i}+\omega_{j}}.$ (1)
The Plackett-Luce model (Plackett, 1975) extended the Bradley-Terry to allow for multiple comparisons and has been justified under Luce’s Choice Axiom (Luce, 1959) and Thurstone’s theory of comparative judgement (Thurstone, 1927; Thompson Jr and Singh, 1967; Yellott Jr, 1977). In this model, a ranking $\pi=\{1\prec 2\prec\dots\prec J\}$ of $J$ objects is assigned probability
$P[\Pi=\pi|\omega_{1},\dots,\omega_{J}]$ $=\prod_{j=1}^{J}\frac{\omega_{j}}{\sum_{j^{\prime}=j}^{J}\omega_{j^{\prime}}},$ (2)
where often one sets $\sum_{j}\omega_{j}=1$ for identifiability. Rankings drawn from the Plackett-Luce model may be interpreted as being created sequentially, where in the first stage an object is selected among all the options, in the second stage an object is selected among all the remaining, and so on. Extensions of distributions in the BTL family have been proposed to capture intricacies in ranked preferences such as order of presentation effects, ties, and covariates (Rao and Kupper, 1967; Critchlow and Fligner, 1991; Gormley and Murphy, 2010; Chapman and Staelin, 1982).
Since BTL distributions have continuous parameters, rank-clusters may be estimated by employing parameter fusion or shrinkage. Parameter fusion is the process of simultaneously estimating parameter values and groups of parameters that should be set equal in value (i.e., “fusing” parameters together). Masarotto and Varin (2012) analyze pairwise comparison data from sports tournaments with parameter fusion techniques under the Bradley-Terry model. Masarotto and Varin (2012) estimate an overall ranking of teams with rank-clusters by applying the frequentist fused lasso (Tibshirani et al., 2005), in which the absolute difference between every pair of worth parameters is penalized after some data-driven normalization. In this approach, the fused parameters are made equal and thus create a rank-cluster among the corresponding objects. The approach of Masarotto and Varin (2012) was extended to additional datasets in sports (Tutz and Schauberger, 2015) and academic journal rankings (Varin et al., 2016; Vana et al., 2016). Jeon and Choi (2018) argued that shrinkage methods like those proposed by Masarotto and Varin (2012) and Tutz and Schauberger (2015) were developed specifically for pairwise comparisons, and thus have inappropriate penalty functions for application to richer kinds of ordinal data like partial or complete rankings. As a result, Jeon and Choi (2018) proposed a modified regularization penalty that may be applied to partial or complete rankings under the Plackett-Luce model.
Parameter fusion methods for rank-clustering exhibit five distinct disadvantages: First, maximum likelihood estimation of models in the BTL family, even in their simplest forms, often suffers from numerical instability and slow computational speed. As a result, numerous authors have proposed complex algorithms to improve estimation accuracy or speed (Hunter et al., 2004; Maystre and Grossglauser, 2015; Turner et al., 2020; Nguyen and Zhang, 2023). Second, uncertainty quantification is challenging and theoretically tenuous in lasso-based methods (Tibshirani, 1996; Fan and Li, 2001). Third, lasso penalty parameters may be difficult to select, requiring data-driven or ad hoc techniques (Tibshirani, 1996; Masarotto and Varin, 2012). Thus, interpretation of the resulting parameter estimates and associated uncertainty is reliant on the specific choice of penalty parameter. Fourth, prior knowledge on the amount and size of rank-clusters cannot be directly incorporated into the frequentist framework: Although the penalty parameter influences estimation of rank-clusters, the specific meaning of various possible choices is not directly interpretable. Fifth, to our knowledge only Jeon and Choi (2018) have formulated a parameter fusion-based method for ordinal data types beyond simple pairwise comparisons. As a result, only their parameter fusion method is available to estimate rank-clusters from complete or partial ranking data.
Many of these disadvantages may be addressed using Bayesian methods. For example, Bayesian counterparts of the lasso and fused lasso—the Bayesian lasso (Park and Casella, 2008) and Bayesian fused lasso (Casella et al., 2010)—allow for valid uncertainty quantification and incorporation of prior knowledge via the selection of interpretable hyperparameters, as opposed to opaque penalty parameters. Bayesian methods for shrinkage and variable selection can be broadly categorized into two classes, continuous shrinkage and spike-and-slab priors. Examples of continuous shrinkage priors include the Bayesian Lasso (i.e., Laplace; Park and Casella (2008)), Bayesian Fused Lasso (Casella et al., 2010), t-distribution (Song and Cheng, 2020), Normal-Exponential-Gamma (Griffin and Brown, 2005; Shimamura et al., 2019), Normal-Gamma (Griffin and Brown, 2010), horseshoe (Carvalho et al., 2010), and Dirichlet-Laplace (Bhattacharya et al., 2015). No continuous shrinkage priors place positive probability on coefficients (or their differences) being precisely zero, which would encourage sparsity. Thus, parameter fusion must be performed via thresholding the posterior distribution, which is often ad-hoc (Porwal and Rodriguez, 2021). This leads us to consider priors of the latter class. Spike-and-slab priors (Mitchell and Beauchamp, 1988; George and McCulloch, 1997; Ishwaran and Rao, 2005) assign weight to both a point-mass at 0 (“spike”) and a continuous density function (“slab”). Although the specific formulations of these priors vary, they allow us to estimate which parameters are precisely zero in a probabilistic framework. Thus, spike-and-slab priors can be used for parameter fusion and uncertainty quantification. However, we are aware of only one existing modification of this prior class for parameter fusion: Wu et al. (2021) apply spike-and-slab to the differences in successive parameters in the linear regression setting. In their method, the order of parameters from least to greatest in coefficient value must be known in advance (as in the fused lasso). This is not practical in the ordinal data setting because the parameter order is equivalent to the overall ranking, whose estimation is a primary goal in the canonical preference data problem. Thus, no Bayesian parameter fusions methods exist which may be directly applied to ordinal data analyses with rank-clustering.
2.2 Methods based on Mallows Distributions
Alternatively, one may consider rank-clustering under the Mallows family of ranking models (Mallows, 1957). The Mallows family is parameterized by the overall ranking, $\pi_{0}$, and a scale parameter $\theta\geq 0$ that dictates how likely rankings of a given distance to $\pi_{0}$ are to be drawn. Specifically, the probability of drawing a ranking $\pi$ from a Mallows$(\pi_{0},\theta)$ distribution is
$P[\Pi=\pi|\pi_{0},\theta]$ $=\frac{e^{-\theta d(\pi,\pi_{0})}}{\psi(\theta)}$ (3)
where $d(\cdot,\cdot)$ is a distance metric and $\psi(\theta)$ is a function which provides an appropriate normalizing constant. Foundational models in the family are defined by their distance metric, with common choices being the Kendall’s $\tau$ (Kendall, 1938) and Spearman’s $\rho$ (Spearman, 1904).
To our knowledge, the Clustered Mallows Model proposed by Piancastelli and Friel (2024) is the only rank-clustering method based on the Mallows model. Their work, proposed concurrently and independently to ours, models item indifference (i.e., rank-clusters) by permitting the overall ranking parameter $\pi_{0}$ to include groups of objects that are tied in rank. The model is estimated in a Bayesian framework from the observed ranking data. However, there are 3 major limitations to their work: First and most importantly, the model requires both the number of rank-clusters and the number of objects per cluster to be pre-specified. Although the authors propose sensible and efficient tools for model selection, the requirement opens the possibility of model misspecification. For example, given 7 objects there are 127 model specifications; given 10 objects there are 1023 model specifications. In addition, pre-specifying the rank-clustering structure removes any uncertainty in the number of rank-clusters and their sizes from the inference task, which we believe to be of key interest in many applications. Second, Bayesian inference of a Clustered Mallows Model is in the class of doubly-intractable problems since the proposed model’s normalizing constant is not available in closed form. As a result, exact inference may be computationally slow, or approximation methods may need to be used that require an inexact pseudolikelihood approach. Third, the Mallows model is best suited for ordinal data in the form of complete or partial rankings, meaning the Clustered Mallows Model cannot handle pairwise or groupwise comparisons. As will be shown in Section 3, our proposed model avoids all three issues by incorporating parameter fusion in the continuously-parameterized Bradley-Terry-Luce model family.
3 Methods
In this section, we first develop a novel spike-and-slab prior for parameter fusion based on partitions. Then, we employ the prior in a model for rank-clustering based on the Bradley-Terry-Luce family of ordinal data models.
3.1 Partition-based Spike-and-Slab Fusion (PSSF) Prior
Suppose data are drawn exchangeably from a model, $\mathcal{M}$, parameterized by the vector $\omega$. We suppose $\omega$ is of length $J$ and let each $\omega_{j}\in\Omega$, $\Omega\subseteq\mathbb{R}$. Our goal is to estimate $\omega$ under the belief that some pairs or groups of parameters in $\omega$ may be clustered (i.e., fused). We say that two parameters $m,n\in\{1,\dots,J\}$, $m\neq n$, are clustered precisely when $\omega_{m}=\omega_{n}$. Clustered parameters may take on any value in their domain, $\Omega$.
Before specifying the prior, we provide some notation on partitions. A partition of an object set $\mathcal{J}=\{1,2,\dots,J\}$ is a collection $g=\{C(1),C(2),\dots,C(K)\}$ of $K$ disjoint nonempty subsets (henceforth referred to as “clusters”) of $\mathcal{J}$ such that their union forms $\mathcal{J}$. Let $C^{-1}(j)$ represent the cluster that contains object $j\in\mathcal{J}$. We let $S(k)=\big{|}\{C(k)\}\big{|}$ be the size of the subset $C(k)$, and denote by $K$ the number of clusters in $g$. To emphasize dependence on $g$, we often write $K_{g}$, $C_{g}(k)$, etc. Lastly, we let $\mathcal{G}$ represent the collection of all partitions $g$ of $\mathcal{J}$, and let $\mathcal{G}_{k}=\{g\in\mathcal{G}:K_{g}=k\}$.
We are now ready to specify the Partition-based Spike-and-Slab Fusion (PSSF) prior. Under PSSF, $\omega$ is assumed to be generated via the following hierarchical model:
$G$ $\sim f_{G}$
$\nu_{k}|G=g$ $\overset{iid}{\sim}f_{\nu}$ $k=1,2,\dots,K_{g}$ (4)
$\omega_{j}$ $=\nu_{C^{-1}_{g}(j)}$ $j\in\mathcal{J}$
In Equation 4, $f_{G}(\cdot)$ is a probability mass function on $\mathcal{G}$ and $f_{\nu}(\cdot)$ is a probability density function on $\Omega$. In words, the prior generates a partition $g$, and then assigns a unique value $\nu_{k}$ to each cluster $C(k)\in g$. Last, each parameter in $\omega$ is assigned the value of $\nu$ corresponding to its cluster in $g$.
As an example, suppose $\mathcal{J}=\{1,2,3\}$ and we draw $g=\{C(1),C(2)\}$ such that $C(1)=\{2\}$ and $C(2)=\{1,3\}$, and draw $\nu=[5,10]$. Then, $\omega=[10,5,10]$ because,
$\omega_{1}$ $=\nu_{C^{-1}_{g}(1)}=\nu_{2}=10,$
$\omega_{2}$ $=\nu_{C^{-1}_{g}(2)}=\nu_{1}=5,\text{ and}$
$\omega_{3}$ $=\nu_{C^{-1}_{g}(3)}=\nu_{2}=10.$
3.1.1 Marginal Prior Probabilities
A useful feature of the PSSF prior is that, regardless of $f_{G}$, the marginal distribution of each $\omega_{j}$ follows $f_{\nu}$. This is because,
$P[\omega_{j}]$ $=\sum_{k=1}^{J}P[\nu_{k}|j\in C(k)]P[j\in C(k)]$ (5)
$=P[\nu_{1}]\sum_{k=1}^{J}P[j\in C(k)]$ (6)
$=f_{\nu}(\cdot).$ (7)
Equation 5 holds as there cannot be more than $J$ clusters and each object belongs to precisely one cluster, Equation 6 holds by the exchangeability of $\nu_{k}$, and Equation 7 holds since $P[\nu_{1}]=f_{\nu}(\cdot)$ by definition and the Law of Total Probability.
3.1.2 Relationship to Spike-and-Slab
We have not yet explained the proposed PSSF prior’s relationship to the spike-and-slab. It is easiest to understand their connection by considering the joint prior distribution on two arbitrary component parameters, $\omega_{m}$ and $\omega_{n}$, such that $m\neq n$. Due to the partitioning structure of parameters in the PSSF prior, there is prior probability associated with a parameter cluster. Thus, their joint prior distribution contains a “spike” component along the line $\omega_{m}=\omega_{n}$, with density of that line determined by $f_{\nu}$. Oppositely, given $\omega_{m}\neq\omega_{n}$ their joint prior distribution reflects independent draws from $f_{\nu}$.
Figure 1 gives examples of the PSSF prior under varying choices of $f_{G}$ and $f_{\nu}$. In all panels, we let $\mathcal{J}=\{1,2\}$ and display the joint prior distribution of $(\omega_{1},\omega_{2})$. In this setting, there are only two unique partitions, $g=\{1,1\}$ and $g=\{1,2\}$. Thus, we specify the prior $f_{G}$ by stating the so-called “cluster probability,” i.e., the probability that $g=\{1,1\}$. Columns correspond to cluster probabilities $0.1,0.5$, and $0.9$, respectively. Rows correspond to $f_{\nu}=\text{Normal}(0,1)$ and $\text{Gamma}(5,3)$, respectively. We notice that as the cluster probability increases, so does the density of points in the spike component. Regardless of $f_{G}$, marginal distributions of each parameter follow $f_{\nu}$. The marginal relationships seen in Figure 1 hold identically even as $\mathcal{J}$ grows.
Additionally, we display the difference between parameters, $\omega_{2}-\omega_{1}$, across different scenarios in Figure 2. The rows and columns are identical to that from Figure 1 and make clear the PSSF prior’s relationship with the traditional spike-and-slab, which has a spike component at 0 and a background slab density.
Figure 1: Joint distribution of $(\omega_{1},\omega_{2})$ under the PSSF prior with varying combinations of $f_{G}$ and $f_{\nu}$. In all cases, $\mathcal{J}=\{1,2\}$, and plots show 20,000 sampled values with marginal density estimates along the axes. Rows correspond to the choice of $f_{\nu}$ and columns to $f_{G}$.
Figure 2: Distribution of $\omega_{2}-\omega_{1}$ under the PSSF prior with varying combinations of $f_{G}$ and $f_{\nu}$. In all cases, $\mathcal{J}=\{1,2\}$. Rows correspond to the choice of $f_{\nu}$ and columns to $f_{G}$.
3.2 Rank-Clustered Bradley-Terry-Luce Model
We now introduce the Rank-Clustered Bradley-Terry-Luce model for ordinal data. Let $I$ be the number of judges who assess $J$ objects. Let $\Pi_{i}$ represent the ordinal preferences provided by judge $i$, which may be a partial ranking, complete ranking, pairwise comparison, or groupwise comparison. Let $R_{i}=|\Pi_{i}|$ and $\mathcal{S}_{i}$ denote the objects in $\mathcal{J}$ considered by judge $i$, such that $\mathcal{S}_{i}\subseteq\mathcal{J}$. $R_{i}$ and $\mathcal{S}_{i}$ are assumed known.
Under the Rank-Clustered Bradley-Terry-Luce (BTL) model, we assume ordinal data is generated via the following Bayesian model:
$\begin{split}\omega&\sim\text{PSSF}(f_{G}\propto\text{Poisson}(K_{g}|\lambda),f_{\nu}=\text{Gamma}(\nu_{k}|a_{\gamma},b_{\gamma}))\\ \Pi_{i}|\omega&\overset{iid}{\sim}\text{Bradley-Terry-Luce}(\omega|\mathcal{S}_{i},R_{i})\hskip 128.0374pti=1,\dots,I\end{split}$ (8)
Rank-Clustered BTL applies the proposed PSSF prior under specific choices of $f_{G}$ and $f\nu$ to the BTL family of distributions for ordinal data. Note that the data-generating BTL distribution is identifiable up to scalar multiplication of $\omega$. However, the proposed Bayesian model does not suffer from identifiability issues due to the non-uniform prior on $\omega$ (Johnson et al., 2022). We emphasize that the model does not pre-specify the number of clusters, a specific rank-clustering structure, or the order of objects. These are treated as random variables and estimated simultaneously.
3.2.1 Prior Selection
We now discuss the selection of priors and hyperparameters. We set $f_{G}$ according to
$f_{G}(g)\propto\text{Poisson}(K_{g}|\lambda).$ (9)
In words, the prior probability of drawing a specific partition $g$ depends only on how many unique clusters, $K_{g}$, it contains. Thus, every partition with the same $K_{g}$ has equal prior probability. As a consequence, cluster sizes do not explicitly impact the prior probability of each $g$. Still, there is an implicit connection between cluster size and $K_{g}$. For example, if $K_{g}=J$, every cluster must be a singleton. In this setup, one could set $\lambda\approx 1$ to encourage rank-clustering, or $\lambda\approx J$ to discourage rank-clustering. Next, we set $f_{\nu}$ according to
$f_{\nu}(\nu_{k})=\text{Gamma}(\nu_{k}|a_{\gamma},b_{\gamma}).$ (10)
This Gamma prior has been used in Bayesian estimation of BTL models as it allows for closed-form Gibbs sampling via data augmentation (Caron and Doucet, 2012; Mollica and Tardella, 2017). The hyperparameters $a_{\gamma}$ and $b_{\gamma}$ control the prior distribution on the worth parameters. Since $\omega$ is invariant to multiplicative transformations, $a_{\gamma}$ and $b_{\gamma}$ are generally non-influential. Nonetheless, because the ratios between worth parameters could become very large when one object is strongly preferred over another, $(a_{\gamma},b_{\gamma})$ should be chosen to give some density to values near 0 to allow for such extreme ratios.
4 Bayesian Estimation
In this section, we develop a Gibbs sampler for Bayesian estimation of Rank-Clustered BTL models and provide simulations to demonstrate its performance under varying numbers of observations and rank-clusters.
4.1 Gibbs Sampler
Equation 4 defines $\omega$ by the pair $(\nu,g)$. Thus, to estimate $\omega$, we sample from the joint posterior distribution of $(\nu,g)$. We do so using a reversible jump Markov chain Monte Carlo (RJMCMC) Gibbs sampler that alternates between updating $g$ and $\nu$ via their full conditionals after data augmentation. The sampler is summarized in Algorithm 1.
Algorithm 1 Gibbs sampler for Rank-Clustered Bradley-Terry-Luce models
  1. 1.
    Initialize $g^{(0)},\nu^{(0)}$ at random, ensuring that $|\nu^{(0)}|=K_{g^{(0)}}$.
  2. 2.
    For $t=1,2,\dots,T_{1}$,
    1. (a)
      Sample $g^{(t)}$ via its full conditional using RJMCMC in order to traverse the space of partitions of varying numbers of clusters.
    2. (b)
      Sample $\nu^{(t)}$ via its full conditional $T_{2}$ times, which is possible via closed-form Gibbs sampling with data augmentation.
Based on our experience fitting Rank-Clustered BTL models to real and simulated data, we recommend initializing $g^{(0)}=\{1,2,\dots,J\}$ (and thus $K_{g^{(0)}}=J$) as it allows rank-clusters to be formed during the estimation process (as opposed to being imposed by the analyst during initialization). For Step 2, $T_{1}$ should be sufficiently large to allow for convergence of the MCMC chain, although specific choices are context-dependent. Step 2(a) performs RJMCMC on clusters of objects. Since RJMCMC can be slow to converge in high dimensions, it is important to run multiple chains and assess for mixing and convergence (Gelman et al., 2013). Step 2(b) relies on a closed-form Gibbs sampler. We find $T_{2}\leq 5$ is usually sufficient for posterior sampling.
4.1.1 Details of Step 2(a)
We now detail Step 2(a), which proposes a new partition $g^{\prime}$ based on the current partition $g$. Since $(g,\nu)$ are intricately tied, $\nu$ must simultaneously be updated to an appropriate $\nu^{\prime}$. The sampling of discrete partitions is challenging to perform efficiently. In a seminal paper on RJMCMC, Green (1995) provided a method for sampling partitions. We adapt that work for the Rank-Clustered BTL model.
Following Green (1995), we only propose $g^{\prime}$ which are slight modifications of $g$: Precisely, we allow only for ‘births’ splitting one cluster into two, or ‘deaths’ merging two clusters into one. Since all partitions have positive probability, this process is irreducible, as required. There is no need to propose $g^{\prime}$ that shuffle the partitions but maintain the number of clusters, as these partitions may be obtained by successive birth and death moves.
[…]
Births and deaths require updating $\nu$ by increasing or decreasing its dimension by 1, respectively. In a birth, we split a cluster’s worth $\nu_{k}$ into $(\nu^{\prime}_{k_{1}},\nu^{\prime}_{k_{2}})$ using,
$\nu^{\prime}_{k_{1}}=u\nu_{k},\ \ \ \nu^{\prime}_{k_{2}}=u^{-1}\nu_{k},$ (11)
where $u\sim\text{Unif}(0.5,1.5)$. The corresponding death solves these equations simultaneously:
$\nu_{k}=\sqrt{\nu^{\prime}_{k_{1}}\nu^{\prime}_{k_{2}}}.$ (12)
For reversibility, we automatically reject proposed births where $\nu^{\prime}_{k_{1}},\nu^{\prime}_{k_{2}}$ are not adjacent.
Per Green (1995), the Metropolis-Hastings probabilities for a birth and death, respectively, are $\min(1,A)$ and $\min(1,A^{-1})$, where
$A$ $=\frac{P(\nu^{\prime},g^{\prime}|\Pi)}{P(\nu,g|\Pi)}\times\frac{q(\nu,g|\nu^{\prime},g^{\prime})}{q(\nu^{\prime},g^{\prime}|\nu,g)P(u)}\times\Big{|}\frac{\partial(\nu^{\prime}_{k_{1}},\nu^{\prime}_{k_{2}})}{\partial(u,\nu_{k})}\Big{|},$ (13)
where $q(\nu^{\prime},g^{\prime}|\nu,g)$ is the transition probability of sampling $(\nu^{\prime},g^{\prime})$ given current parameter set $(\nu,g)$. We now calculate each term in $A$. First,
[…]
$=\frac{P(\Pi|\nu^{\prime},g^{\prime})P[\nu^{\prime}|g^{\prime}]P[g^{\prime}]}{P(\Pi|\nu,g)P[\nu|g]P[g]}$
$=\frac{P(\Pi|\nu^{\prime},g^{\prime})}{P(\Pi|\nu,g)}\times\frac{\text{Gamma}(\nu^{\prime}_{k_{1}}|a_{\gamma},b_{\gamma})\text{Gamma}(\nu^{\prime}_{k_{2}}|a_{\gamma},b_{\gamma})}{\text{Gamma}(\nu_{k}|a_{\gamma},b_{\gamma})}\times\frac{P[g^{\prime}]}{P[g]},$ (14)
where $P(\Pi|\nu,g)$ and $P[g]$ are defined by Equation 8. Second,
$\frac{q(\nu,g|\nu^{\prime},g^{\prime})}{q(\nu^{\prime},g^{\prime}|\nu,g)P(u)}$ $=\frac{d_{g^{\prime}}\times\frac{1}{K_{g^{\prime}}-1}}{\Big{(}b_{g}\times\frac{1}{\#\{l:S_{l}(g)\geq 2\}}\times\frac{2}{2^{S_{g}(k)}-2}\Big{)}\Big{(}\frac{1}{1.5-0.5}\Big{)}}$ (15)
$=\frac{d_{g^{\prime}}\#\{l:S_{g}(l)\geq 2\}(2^{S_{g}(k)-1}-1)}{b_{g}(K_{g^{\prime}}-1)}$
The numerator in Equation 15 is the death probability, $d_{g^{\prime}}$, times the probability of selecting a pair of adjacent partitions given $K_{g^{\prime}}$ total partitions after a split (there are $K_{g^{\prime}}-1$ such pairs). The denominator is the birth probability, $b_{g}$, times the probability of selecting a specific cluster $k$ among those with at least two members. This term also includes the probability of dividing the $S_{g}(k)$ objects in cluster $k$ into two non-empty subsets. There are $(2^{S_{g}(k)}-2)/2$ such subsets, since there are $2^{S_{g}(k)}$ total possible partitions, two empty partitions, and two ways to obtain each two-way split. Third and last,
$\Big{|}\frac{\partial(\nu^{\prime}_{k_{1}},\nu^{\prime}_{k_{2}})}{\partial(u,\nu_{k})}\Big{|}$ $=\Bigg{|}\begin{bmatrix}\frac{\partial}{\partial u}\nu_{k_{1}}^{\prime}&\frac{\partial}{\partial\nu_{k}}\nu_{k_{1}}^{\prime}\\ \frac{\partial}{\partial u}\nu_{k_{2}}^{\prime}&\frac{\partial}{\partial\nu_{k}}\nu_{k_{2}}^{\prime}\end{bmatrix}\Bigg{|}=\Bigg{|}\begin{bmatrix}\frac{\partial}{\partial u}u\nu_{k}&\frac{\partial}{\partial\nu_{k}}u\nu_{k}\\ \frac{\partial}{\partial u}\nu_{k}/u&\frac{\partial}{\partial\nu_{k}}\nu_{k}/u\end{bmatrix}\Bigg{|}=\Bigg{|}\begin{bmatrix}\nu_{k}&u\\ -\nu_{k}/u^{2}&1/u\end{bmatrix}\Bigg{|}$
$=\frac{2\nu_{k}}{u}.$ (16)
4.1.2 Details of Step 2(b)
To update $\nu$ conditional on a partition $g$ and our data, $\Pi$, we turn to a clever data augmentation trick for Bayesian estimation of Plackett-Luce models as seen in Caron and Doucet (2012) and Mollica and Tardella (2017). Here, we adapt their trick to account for the more general BTL family of distributions and rank-clustering. Let $Y=\{Y_{ir}\}$ be a collection of independent random variables, $i=1,\dots,I$ and $r=1,\dots,R_{i}$, sampled according to
$Y_{ir}\sim\text{Exponential}\Big{(}\sum_{j\in\mathcal{S}_{i}}\nu_{g^{-1}(j)}-\sum_{s=0}^{r-1}\nu_{g^{-1}(\pi_{i}(s))}\Big{)}.$ (17)
The exponential rates are precisely the denominator terms from BTL densities that are burdensome to calculate. The full conditional posterior probability $P[\nu|Y,\Pi,g]$ is then,
$P[\nu|Y,\Pi,g]\propto$ $P[Y|\Pi,g,\nu]P[\Pi|g,\nu]P[g|\nu]P[\nu]$
$\propto$ $P[Y|\Pi,g,\nu]P[\Pi|g,\nu]P[\nu]$
$=$ $\prod_{i=1}^{I}\prod_{r=1}^{R_{i}}\Big{(}\sum_{j\in\mathcal{S}_{i}}\nu_{g^{-1}(j)}-\sum_{s=0}^{r-1}\nu_{g^{-1}(\pi_{i}(s))}\Big{)}e^{-y_{ir}\big{(}\sum_{j\in\mathcal{S}_{i}}\nu_{g^{-1}(j)}-\sum_{s=0}^{r-1}\nu_{g^{-1}(\pi_{i}(s))}\big{)}}\times$
$\prod_{i=1}^{I}\prod_{r=1}^{R_{i}}\frac{\nu_{g^{-1}(\pi_{i}(r))}}{\sum_{j\in\mathcal{S}_{i}}\nu_{g^{-1}(j)}-\sum_{s=0}^{r-1}\nu_{g^{-1}(\pi_{i}(s))}}\times\prod_{k=1}^{K}\nu_{k}^{a_{\gamma}-1}e^{-b_{\gamma}\nu_{k}}$
$=$ $\prod_{i=1}^{I}\prod_{r=1}^{R_{i}}\nu_{g^{-1}(\pi_{i}(r))}e^{-y_{ir}\big{(}\sum_{j\in\mathcal{S}_{i}}\nu_{g^{-1}(j)}-\sum_{s=0}^{r-1}\nu_{g^{-1}(\pi_{i}(s))}\big{)}}\times\prod_{k=1}^{K}\nu_{k}^{a_{\gamma}-1}e^{-b_{\gamma}\nu_{k}}$ (18)
Given these cancellations, we notice a closed-form expression for the posterior:
$P[\nu|Y,\Pi,g]\propto$ $\prod_{i=1}^{I}\prod_{k=1}^{K}\nu_{k}^{c_{ki}}e^{-\nu_{k}\sum_{r=1}^{R_{i}}y_{ir}\delta_{irk}}\times\prod_{k=1}^{K}\nu_{k}^{a_{\gamma}-1}e^{-b_{\gamma}\nu_{k}}$
$=$ $\prod_{k=1}^{K}\nu_{k}^{a_{\gamma}+\sum\limits_{i=1}^{I}c_{ki}-1}e^{-\nu_{k}(b_{\gamma}+\sum\limits_{i=1}^{I}\sum\limits_{r=1}^{R_{i}}y_{ir}\delta_{irk})}$
$\propto$ $\prod_{k=1}^{K}\text{Gamma}\Big{(}\nu_{k}\ \Big{|}\ a_{\gamma}+\sum\limits_{i=1}^{I}c_{ki},b_{\gamma}+\sum\limits_{i=1}^{I}\sum\limits_{r=1}^{R_{i}}y_{ir}\delta_{irk}\Big{)}$ (19)
where
$c_{ki}$ =|$=\big{|}\big{\{}j:j\in\pi_{i},g^{-1}(j)=k\big{\}}\big{|}$ (20)
$\delta_{irk}$ =|$=\big{|}\big{\{}j:j\in\mathcal{S}_{i},j\not\in\{\pi_{i}(1),\dots,\pi_{i}(r-1)\},g^{-1}(j)=k\big{\}}\big{|}.$ (21)
Thus, we can sample $\nu$ from a closed-form Gamma distribution after augmentation of the conditioning data $\Pi$ and random variable $g$ with $Y$.
Now that we have developed an efficient estimation algorithm for Rank-Clustered BTL models, we turn to a numerical simulation to demonstrate estimation accuracy under different rank-clustering regimes.
4.2 Numerical Simulation
We now demonstrate accurate estimation of worth parameters and rank-clusters via a Rank-Clustered BTL model in a numerical simulation. We assume there are $J=8$ objects which form $K$=1, 2, 4, or 8 rank-clusters. When $K=J=8$, every object is independent; there are only singleton rank-clusters. In the true worth parameter vector, $\omega_{0}$, rank-clustered objects have identical values and successive rank-clusters are separated in value by a factor of 4 (see Table 1 for specific values). Fourfold increases induce strong but not absolute separation between objects: For demonstration, in a pairwise tournament between an object with $\omega_{1}=1$ and $\omega_{2}=4$, the probability of selecting object 2 is,
$P[2\prec 1|\omega_{1}=1,\omega_{2}=4]=\frac{\omega_{2}}{\omega_{1}+\omega_{2}}=\frac{4}{4+1}=0.8.$
We also vary the Poisson hyperparameter on the number of rank-clusters, $\lambda\in\{0.1,2,4,8\}$, which encourages rank-clustering to different extents and allows us to measure robustness of results when $\lambda$ is somewhat misspecified. Finally, we vary the number of judges $I\in\{50,200,800\}$. For each combination of $I$, $\omega_{0}$, and $\lambda$, we generate 20 independent datasets and fit a Rank-Clustered Bradley-Terry-Luce distribution to each, under hyperparameters $a_{\gamma}=5$ and $b_{\gamma}=3$. We set $T_{1}=5{,}000$ and $T_{2}=2$ to obtain 10,000 posterior samples in each MCMC chain and remove the first half as burn-in. For identifiability, posterior estimates of $\omega_{0}$ are normalized post-hoc such that $\sum_{j}\omega_{0j}=1$.
Setting: $\omega_{0}$
$K=1$$\{4^{0},4^{0},4^{0},4^{0},4^{0},4^{0},4^{0},4^{0}\}$
[…]
$K=8$$\{4^{0},4^{1},4^{2},4^{3},4^{4},4^{5},4^{6},4^{7}\}$
Table 1: Simulation settings for $\omega_{0}$ under varying numbers of true rank-clusters, $K$.
We first examine the accuracy of estimation for $\omega_{0}$ across simulation settings. Figure 3 displays boxplots of mean absolute error (MAE) for $\omega_{0}$ by number of judges $I$, true number of rank-clusters $K$, and the choice of hyperparameter $\lambda$. In general, estimation is quite accurate. We see that for any specific combination of $K$ and $\lambda$, MAE decreases as $I$ increases. Estimation error is higher when $K$ is large and $I$ is small, most likely the result of error estimating a complex rank-clustering structure.
Figure 3: Boxplots of posterior mean absolute error for $\omega_{0}$ across combinations of the number of judges $I$, true number of rank-clusters $K$, and hyperparameter $\lambda$. Errors are calculated after normalization of posterior samples such that $\sum_{j}\omega_{0j}=1$.
Figure 4 displays the mean posterior probability of rank-clustering across object pairs which are truly rank-clustered (navy) or independent (gold) in $\omega_{0}$.
Figure 4: Boxplots of the mean posterior probability of rank-clustering object pairs which are truly rank-clustered (navy) or independent (gold), across combinations of $I$, $K$, and $\lambda$.
Results are further separated by the number of judges, $I$, true number of clusters, $K$, and hyperparameter $\lambda$. For rank-clustered pairs, accuracy of recovery is generally high and increases with the number of judges, $I$. Accuracy is best when hyperparameter $\lambda\approx K$, which occurs when prior belief regarding the number of rank-clusters is approximately correct. If there is limited prior knowledge on the number of rank-clusters, we suggest specifying a vague hyperparameter setting such as $\lambda=\frac{J}{2}$ and assessing sensitivity of results to various choices of $\lambda$. The posterior probability of rank-clustering independent object pairs is near 0 in all simulations, indicating excellent recovery accuracy of objects with distinct worth parameters.
The numerical simulations in this section indicate that the proposed Rank-Clustered BTL model is able to accurately estimate the relative worth of objects in a collection, including in the presence of rank-clustering. Estimation error decreases to 0 as the number of observations increases. Overall, the model correctly identifies rank-clustered and independent object pairs.
5 Applications
In this section, we apply our model to four real datasets involving ordinal comparisons. These four applications were chosen to highlight the applicability of our method to various ordinal data types and domain areas, which are summarized in Table 2. The data sets are comprised of sushi preferences of Japanese adults (Kamishima, 2003), policy preferences of respondents from Great Britain in a Eurobarometer survey (Reif and Melich, 1993), ranked-choice votes in a Minneapolis mayoral election (Minneapolis Elections and Voter Services, 2021), and pairwise game outcomes among teams in the National Basketball Association (National Basketball Association, 2021).
Section Setting Methodological Value
5.1 Sushi preferences in Tohoku Rank-clustering with complete rankings; comparing inferred overall ranking with that of the Clustered Mallows Model.
5.2 Minneapolis mayoral election votes Rank-clustering with top-3 partial rankings among 17 candidates; interpretable overall ranking captures the winner’s mandate in ranked-choice elections; comparing inferred overall ranking with those from a standard BTL model and two election procedures.
5.3 Eurobarometer survey policy preferences Rank-clustering with partial rankings; inferred overall ranking permits indistinguishability among options to aid policymakers.
5.4 Basketball game outcomes Rank-clustering with pairwise comparison data; inferred overall order of captures similarly-performing teams.
Table 2: Summary of applications.
5.1 Sushi Preferences in Tohoku
We first study complete preference rankings of 10 common sushi types based on a benchmarking dataset by Kamishima (2003). To allow our results to be comparable with an analysis of the sushi data in the related work of Piancastelli and Friel (2024), we analyze the preferences of survey respondents who lived in Japan’s Tohoku region until at least 15 years of age. In total, there were 280 respondents providing complete rankings of 10 sushi types. We fit a Rank-Clustered BTL distribution to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering but permit a wide variety of outcomes. Figure 5 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 5: Primary results of an analysis of preference rankings of 10 common sushi types in a survey of 280 Japanese adults. Posterior rank-clustering probabilities (left) and posterior distributions of sushi-specific worth parameters (right).
Sushi types are ordered according to their posterior median worth. Based on the left panel in Figure 5, fatty tuna appears to be the strictly most preferred sushi type based on this population, followed by tuna and shrimp rank-clustered with high posterior probability in second place. Salmon roe and sea eel exhibit high posterior probability of rank-clustering, as do sea urchin, tuna roll, and squid; these two groups may themselves be rank-clustered. Egg and cucumber roll are rank-clustered in last place. Our results demonstrate the proposed model’s ability to rank-cluster objects with uncertainty under complete rankings in survey data.
[…]
5.2 2021 Minneapolis Mayoral Election
Our second example analyzes real rank-choice votes from the 2021 mayoral election in Minneapolis, Minnesota (Minneapolis Elections and Voter Services, 2021). This election included 17 candidates (excluding write-ins and one who received no votes) and asked voters to rank their top-3 choices, in order. A total of 145,337 votes were cast in this election. To mimic exit polling data, we randomly sample 1000 valid votes for analysis, which we treat as a random sample of preferences from the population of Minneapolis voters. We want to estimate the overall preferences of Minneapolis voters regarding mayoral candidates and learn which candidates, if any, are rank-clustered at the population level. Clustering candidates may be of interest to political scientists or local political organizations for the purpose of understanding voter preferences (Gunther and Diamond, 2003; Dimock et al., 2014). For example, if the winner of the election is deemed to be rank-clustered with other candidate(s), their mandate may be considered weak. Conversely, if the winner is a singleton first-place rank-cluster—clearly ranked above all other candidates—their mandate may be considered strong. We fit a Rank-Clustered BTL to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage few of rank-clusters. Figure 6 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 6: Primary results of an analysis of 1000 valid votes in the 2021 Minneapolis mayoral election. Posterior rank-clustering probabilities (left) and posterior distributions of candidate-specific worth parameters (right).
In Figure 6, candidates are ordered by their posterior median estimate of worth. Cluster 1 consists of Jacob Frey, the winner and incumbent. We note that Frey is not rank-clustered with other candidates with high posterior probability, suggesting a relatively strong mandate. Cluster 2 consists of Kate Knuth and Sheila Nezhad, both non-incumbent DFL candidates with substantial political backing and fundraising. Clusters 3 and 4 consist of AJ Awed and Clint Conner, respectively, who are each DFL candidates with less experience and political backing than candidates in previous rank-clusters. Cluster 5 consists of Laverne Turner, a GOP candidate, and Cluster 6 consists of Carney Jr., Harcus, Globus, and Nelson. Last, Cluster 7 consists of 6 candidates with minimal support.
Figure 7 compares point estimates of rank for each candidate among four methods.
Figure 7: Comparison of estimated rank for each candidate across four aggregation methods: Ranked Choice, First-Past-the-Post (FPP), BTL, and Rank-Clustered BTL (RC BTL). Candidates are ordered by their rank in the actual ranked choice election.
The first and second rows display assigned ranks from ranked choice and “first-past-the-post” (FPP) election procedures, respectively. [^3] Since ranked choice and FPP elections use deterministic procedures, we use the term “assigned rank”. We calculate FPP ranks by ordering candidates by the number of first place votes he/she received (ignoring all second and third place votes). [^4] If the actual election had utilized FPP tabulation, results may have been different based on the differing voter strategies encouraged by ranked choice and FPP elections. The third and fourth rows display estimated ranks from a standard Bayesian BTL and our Rank-Clustered BTL, respectively. Estimated ranks are based on maximum a posteriori estimates. Frey wins the election in all methods. The BTL and Rank-Clustered BTL models roughly reflect the deterministic algorithms, although we notice some swaps in candidate ranks which may be attributed to differences between first place and second or third place votes. For example, Conner received fewer first place votes than Turner, but far more second and third place votes. As a result, deterministic algorithms rank Turner above Conner, while the BTL model takes into account the additional preference information and ranks Conner above Turner. In summary, the overall ordering estimated by the Rank-Clustered BTL differs from a standard BTL model and two deterministic election procedures. Furthermore, our model confirms that Frey is strictly preferred over the remaining candidates by voters.
5.3 Eurobarometer 34.1 Survey Data
We analyze data from the Eurobarometer 34.1 survey (Reif and Melich, 1993), which included the following question:
Question 28: There are various actions that could be taken to eliminate the drugs problem. In your opinion, what is the first priority? And the next most urgent? (Ask respondent to rank all 7, with 1 as the most urgent.)
  1. 1.
    Information campaigns about the dangers of drugs.
  2. 2.
    Hunting down drug pushers and distributors.
  3. 3.
    Legal penalty for drug taking.
  4. 4.
    Looking after and treating drug addicts and rehabilitating them.
  5. 5.
    Funding research into drug substitutes, and into the treatment of drug addiction.
  6. 6.
    Fighting the social causes of drug addiction.
  7. 7.
    Reinforcing the control or distribution and usage of addictive medicines.
We subset the data to respondents from Great Britain to avoid heterogeneity and non-proportional sampling among respondents from different European countries. There were 1005 valid responses among this group (out of 1031 total surveyed), of which 970 were complete rankings and the rest ranked between 1 and 5 items (a top-6 ranking is inherently equivalent to a complete ranking since all survey options were presented). We seek to identify a population-level ordering of the priorities that accounts for potential equality or indistinguishability among the options based on the survey data. We note that these data were previously studied by Wang et al. (2017) in a mixed-membership model. Our analysis, although a simplification of the diverse population’s heterogeneous preferences, provides a simpler interpretation to policy-makers interested in understanding holistic attitudes of their constituent population.
We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering. Figure 8 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 8: Primary results of an analysis of respondents from Great Britain to question 28 in the Eurobarometer 34.1 survey. Posterior rank-clustering probabilities (left) and posterior distributions of survey option-specific parameters (right).
We observe that policy option 2 (information campaigns) is strictly preferred to the rest among the population of survey respondents from Great Britain, whereas options 5 (funding research) and 3 (legal penalty) are rank-clustered last. The results indicates to policymakers that respondents in Great Britain strongly prioritize Option 2 in comparison to the rest, while pairs of Options 1 and 4 and Options 3 and 5, are, respectively, indistinguishable within each pair, with 1 and 4 being strongly preferred to 3 and 5. By rank-clustering similarly-preferred options, interpretation of the population’s preferences is simplified for policymakers.
5.4 2020-21 National Basketball Association Game Outcomes
Finally, we analyze outcomes of 1,249 pairwise games in the 2020-21 season of the National Basketball Association (NBA). In this season, there were 30 teams, who each played every other team at least twice; 3 pairs of teams played 11 matches. We seek to estimate an overall ranking of teams that allows for potential equality in ranking.
We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=0.1$ to encourage rank-clustering given the limited ordinal comparison data provided by pairwise matchups. Figure 9 displays posterior rank-clustering probabilities (left) and parameter posteriors (right). Additional results and convergence diagnostics provided in the supplementary materials.
Figure 9: Posterior rank-clustering probabilities (left) and posterior distributions of basketball teams (right) based on the results of all 2020-21 regular season games.
In this setting, the Rank-Clustered BTL model estimates an ordering of professional basketball teams with uncertain rank-clustering patterns. Note than in the left panel of Figure 9, the squares on the main diagonal are black (indicating probability 1) by default since each team is inherently rank-clustered with itself. An 80% posterior credible interval indicates that there are between 4 and 6 rank-clusters; every team has less than 0.015 probability of belonging to a singleton rank-cluster. However, rank-clusters in this setting are less certain than in previous examples. Most clearly, 5 teams (Detroit Pistons, Oklahoma City Thunder, Cleveland Cavaliers, Orlando Magic, and the Houston Rockets) appear to be rank-clustered for last place. Among these 5 teams, pairwise rank-clustering probabilities are at least 0.75. Instead of ranking these teams with granularity, we may simply infer by our model that they were each approximately the worst teams of the league in this season. Conversely at the top of the overall ranking, we observe that the top 5 teams (Utah Jazz, Phoenix Suns, Philadelphia 76ers, Milwaukee Bucks, and the Brooklyn Nets) all have a high posterior probability of rank-clustering. Although the Milwaukee Bucks were the ultimately this season’s league champions, our model suggests that these 5 teams were of roughly indistinguishable ability based on the game data. These summaries from our model, despite not accounting for the complexities of the sport, provide useful and interpretable summaries of the teams’ abilities across the season.
6 Discussion
In this paper, we proposed the Rank-Clustered Bradley-Terry-Luce model for estimating an overall ranking of objects with rank-clusters. The model employs the Bradley-Terry-Luce (BTL) family of distributions for ordinal comparison data and estimates model parameters in a Bayesian framework under the proposed Partition-based Spike-and-Slab (PSSF) prior. The model requires no pre-specification of the number or size of the rank-clusters (as in Piancastelli and Friel (2024)), nor specification of lasso-based penalty parameters (as in Masarotto and Varin (2012); Jeon and Choi (2018)). In a simulation study, we demonstrated the model’s ability to accurately and consistently estimate the relative worth of objects in a collection while simultaneously estimating rank-clusterings over objects. We used Rank-Clustered BTL on four real datasets under different types of ordinal comparison data.
In contrast to the only other spike-and-slab based prior for parameter fusion Wu et al. (2021), PSSF prior we developed does not require a known parameter order. Visual inspection of the prior distribution makes obvious its connection to spike-and-slab: “spike” components correspond to parameter clusters and “slab” components correspond to independent parameters. Estimation of parameters under this model requires reversible jump MCMC. To overcome potentially slow or computationally-burdensome estimation in this setting, we proposed a computationally efficient Gibbs sampler that alternates between updating the partition of objects (based on the seminal work of Green (1995) and updating object-level worth parameters by adapting a data augmentation trick for standard Plackett-Luce models by Caron and Doucet (2012) and later adapted for Plackett-Luce mixtures by Mollica and Tardella (2017).
A useful benefit of estimating parameter values and clusters in a single Bayesian framework is the avoidance of issues associated with selective inference (Taylor and Tibshirani, 2015) or more colloquially, double dipping (Kriegeskorte et al., 2009). Selective inference occurs when the same data is used twice in the process of model selection and/or estimation, e.g., to estimate some latent structure underlying the data and subsequently to estimate parameters conditional on that estimated structure. In our context, selective inference would occur if ordinal preference data was used first to identify rank-clusters and then used again to estimate worth parameter values conditional on those clusters. We note that selective inference occurs in the estimation of the related Clustered Mallows Model by Piancastelli and Friel (2024), which requires selecting the number and size of rank-clusters among objects before fitting the model. Selective inference often leads to invalid inference in part because uncertainty regarding the estimated clustering structure is not taken into account. However, Rank-Clustered Bradley-Terry-Luce models do not perform selective inference because parameter values and rank-clusters are estimated simultaneously. As such, our parameter estimates incorporate uncertainty across the posterior distributions of both the rank-clustering structure and the specific parameter values.
Results from Rank-Clustered BTL models are useful in a variety of contexts. As noted in other fusion literatures on rankings, estimated overall rankings may be easier to understand and interpret when rank-clusters of objects are identified, as rank-clusters lead to fewer rank levels of objects to distinguish (Masarotto and Varin, 2012). In contexts where model results are used for prediction, such as in sports, estimating rank-clusters may improve predictive accuracy (Tutz and Schauberger, 2015). Similarly, estimating rank-clusters is important in the context of decision-making: In peer review, for example, rank-clusters can be beneficial for communicating uncertainty in the assessment of preferences and for better transparency in funding decisions. We might imagine a scenario where a government agency is only able to fund two grants, however, two grant proposals are rank-clustered in second place. In this case, rank-clustering can be used to communicate uncertainty in the relative quality of the top proposals. A potential danger is that under this uncertainty, decision makers may be tempted to resort to unfair tie-breaking methods, e.g., selecting the proposal with the most famous author. Instead, tie-breaking should occur based on a fairer or more principled method, such as a partial lottery (Fang and Casadevall, 2016; Roumbanis, 2019; Heyard et al., 2022).
We list a few possible directions for future research. First, in this work we have not considered the level of interconnectedness among the assessed objects (e.g., if separate groups of judges assess completely distinct sets of objects). This is particularly relevant in the case of pairwise comparison data, in which some pairs of objects may never experience a head-to-head match-up. Second, the PSSF prior could be imposed as a prior for more complex BTL models or to other models entirely. In the former, the PSSF prior could be applied to preference learning via BTL distributions that incorporate covariates (e.g., Baldassarre et al. (2023); Gormley and Murphy (2010); Chapman and Staelin (1982)). In that case, the prior may be modified to permit covariate parameter estimation in addition to rank-clustering. In the latter case, the PSSF prior may be applied to regression for variable fusion, and its performance may be compared to other existing Bayesian variable fusion methods (e.g., Casella et al. (2010); Song and Cheng (2020); Shimamura et al. (2019)).
The proposed Rank-Clustered BTL model accurately estimates rank-clusters, permitting complex summaries beyond the traditional overall ranking and allowing for improved interpretability of the results. The Bayesian Rank-Clustered BTL model relies on a novel, spike-and-slab type prior for parameter fusion, and is estimated in a computationally-efficient manner. The applications in survey data, voting, and sports to aid informed inference and decision-making illustrate methodological versatility and broad applicability of our proposed rank-clustering approach.
Implementation
Code to implement the Rank-Clustered BTL model may be found at https://github.com/pearce790/rankclust.
Acknowledgements
The authors would like to thank Drs. T. Brendan Murphy and I. Claire Gormley for inspiring conversations on rank data modeling during the early stages of this work.
Funding
The authors were supported by the National Science Foundation under Grant No. 2019901.
Appendix Appendix A Additional Application Results
Appendix A.1 Additional Results from Section 5.1
Figure 10 displays stacked bar charts of ranks received by each sushi type by the survey respondents from Tohoku.
Figure 10: Stacked bar charts of ranks received by each sushi type.
Figures 11 and 12 contain trace plots for $K$ and $\omega$ after burn-in for each chain. We find the trace plots to demonstrate satisfactory mixing and convergence.
Figure 11: Trace plot of K in the Sushi data analysis
Figure 12: Trace plots of $\omega$ in the Sushi data analysis
Appendix A.2 Additional Results from Section 5.2
Figure 13 displays stacked bar charts of the sampled votes by rank level for each candidate.
Figure 13: Number of votes by rank level and candidate. Candidates are ordered by their position in the official ranked choice election. Acronyms on the tops of bars represent each candidate’s political party.
Candidates are ordered by their final placement according to the official ranked choice voting algorithm. The incumbent, Jacob Frey, receives the largest share of first place votes, although Kate Knuth and Sheila Nezhad also receive substantial support. The remaining candidates receive comparatively few votes. Most candidates are associated with the Democratic-Farmer-Labor (DFL) party, which is affiliated with the national Democratic Party. Laverne Turner and Bob “Again” Carney Jr. are the only Republicans (GOP) in the race. The remaining candidates represent Grassroots–Legalize Cannabis (GLC), Libertarian (LIB), Socialist Workers Party (SWP), For the People Party (FPP), Independence (INC), Independent (IND), and Humanitarian–Community Party (HCP).
Figures 14 and 15 contain trace plots for $K$ and $\omega$ after burn-in for each chain. We find the trace plots to demonstrate satisfactory mixing and convergence.
Figure 14: Trace plots of $K$ in the Minneapolis mayoral election data analysis.
Figure 15: Trace plots of $\omega$ in the Minneapolis mayoral election data analysis.
Appendix A.3 Additional Results from Section 5.3
Figure 16 displays stacked bar charts of ranks received by each policy option by the survey respondents from Great Britain.
Figure 16: Stacked bar charts of ranks received by each policy option.
Figures 17 and 18 contain trace plots for $K$ and $\omega$ after burn-in for each chain. We find the trace plots to demonstrate satisfactory mixing and convergence.
Figure 17: Trace plot of K in the Eurobarometer survey data analysis
Figure 18: Trace plot of $\omega$ in the Eurobarometer survey data analysis
Appendix A.4 Additional Results from Section 5.4
Figure 19 displays stacked bar charts of the season record of each NBA team across the 2020-21 season.
Figure 19: Stacked bar charts of ranks received by each NBA team across the 2020-21 season. Winning = rank 1; losing = rank 2.
Figures 20 and 21 contain trace plots for $K$ and $\omega$ after burn-in for each chain. We find the trace plots to demonstrate satisfactory mixing and convergence.
Figure 20: Trace plot of K in the 2020-21 NBA season analysis
Figure 21: Trace plot of $\omega$ in the 2020-21 NBA season analysis

### v2

@@
 Abstract
-In a traditional analysis of ordinal comparison data, the goal is to infer an overall ranking of objects from best to worst with each object having a unique rank.
-However, the ranks of some objects may not be statistically distinguishable.
-This could happen due to insufficient data or to the true underlying abilities or qualities being equal for some objects.
-In such cases, practitioners may prefer an overall ranking where groups of objects are allowed to have equal ranks or to be rank-clustered.
+Traditional statistical inference on ordinal comparison data results in an overall ranking of objects, e.g., from best to worst, with each object having a unique rank.
+However, ranks of some objects may not be statistically distinguishable.
+This could happen due to insufficient data or to the true underlying object qualities being equal.
+Because uncertainty communication in estimates of overall rankings is notoriously difficult, we take a different approach and allow groups of objects to have equal ranks or be rank-clustered in our model.
 Existing models related to rank-clustering are limited by their inability to handle a variety of ordinal data types, to quantify uncertainty, or by the need to pre-specify the number and size of potential rank-clusters.
-We solve these limitations through the proposed Bayesian Rank-Clustered Bradley-Terry-Luce model.
-We allow for rank-clustering via parameter fusion by imposing a novel spike-and-slab prior on object-specific worth parameters in Bradley-Terry-Luce family of distributions for ordinal comparisons.
-We demonstrate the model on simulated and real datasets in survey analysis, elections, and sports.
+We solve these limitations through our proposed Bayesian Rank-Clustered Bradley-Terry-Luce model.
+We accommodate rank-clustering via parameter fusion by imposing a novel spike-and-slab prior on object-specific worth parameters in Bradley-Terry-Luce family of distributions for ordinal comparisons.
+We demonstrate rank-clustering on simulated and real datasets in surveys, elections, and sports analytics.
 Keywords ordinal comparisons  $\cdot$ Plackett-Luce  $\cdot$ Bradley-Terry  $\cdot$ spike-and-slab  $\cdot$ fusion priors  $\cdot$ item indifference
@@
 Each judge’s ordinal preferences, $\Pi_{i}$, may be provided in various forms, such as complete rankings, partial rankings, or pairwise comparisons among the available objects or some subset thereof.
-Standard statistical model families for ranking data (e.g., Mallows (Mallows, 1957), Bradley-Terry-Luce (Bradley and Terry, 1952; Plackett, 1975; Luce, 1959)) derive or estimate the rank of each object such that each object receives a unique rank.
-An overall ranking orders all objects from best to worst on the basis of their estimated rank.
-Analyses of this kind are used to rank candidates in an election given ranked choice votes (Gormley and Murphy, 2008; Mollica and Tardella, 2017), sports teams or players in a league using game outcomes (Tutz and Schauberger, 2015; Barrientos et al., 2023), or genes based on ordinal comparisons of genomic data (Eliseussen et al., 2023; Vitelli et al., 2018).
+Standard statistical model families for ranking data such as Mallows (Mallows, 1957) or Bradley-Terry-Luce (Bradley and Terry, 1952; Plackett, 1975; Luce, 1959) derive or estimate the rank of each object whereby each object receives a unique rank.
+An estimated overall ranking then orders all objects from best to worst.
+Analyses of this kind are used to rank candidates in ranked choice elections, (Gormley and Murphy, 2008; Mollica and Tardella, 2017), sports teams or players in a league using pairwise game outcomes (Tutz and Schauberger, 2015; Barrientos et al., 2023), or genes based on ordinal comparisons of genomics data (Eliseussen et al., 2023; Vitelli et al., 2018).
 However, requiring estimated ranks to be unique is not always useful or appropriate.
-Specifically, some objects may be equal or indistinguishable in quality or ability at the population level.
-In such cases, we refer to the objects as being rank-clustered.
-For example, consider one situation of an election in which 2 candidates, both of the same political party, are running for an office.
-If voters express their preferences solely on the basis of party, the candidates are inherently equal in quality; the rational voter would randomly select a candidate to vote for.
-In another case, when the number of votes cast is small, estimated ranks assigned to each candidate could exhibit substantial uncertainty, suggesting the candidates are indistinguishable in quality based on the limited number of observed votes.
-In such situations, allowing for rank-clusters may improve interpretability and aid accurate inference, prediction, and decision-making when analyzing ordinal preferences.
-In this paper, we propose a Bayesian framework for ordinal data analysis that estimates an overall ranking of objects with rank-clusters.
+For example, some objects may be equal or indistinguishable in their true quality or ability.
+Consider an election in which 2 candidates, both of the same political party, are running for an office.
+If voters express their preferences solely on the basis of party, the candidates are inherently equal in quality.
+In another situation, when the number of votes cast is small, estimated ranks assigned to each candidate could exhibit substantial uncertainty, suggesting the candidates are indistinguishable in quality based on the limited number of observed votes.
+In such situations, allowing for inference to estimate the candidates as having the same rank or be rank-clustered may improve interpretability, prediction, and decision-making when analyzing ordinal preferences.
+In this paper, we propose a Bayesian framework for ordinal data analysis that estimates an overall ranking of objects with rank-clusters, develop a computationally-efficient Gibbs sampler for estimation, and apply the model to real and simulated data.
 Specifically, we choose to model observed rank via the Bradley-Terry-Luce (BTL) family of distributions which permits analysis of ordinal preferences in many forms, such as complete rankings, partial rankings, pairwise comparisons, and groupwise comparisons.
-To induce rank clusters, we place a novel spike-and-slab prior that partitions continuous parameters, proposed herein, on the object-specific parameters of BTL distributions.
-In contrast to existing work related to rank-clustering in the literature, our model does not require the parameter order nor the number or size of rank-clusters to be known in advance.
-Instead, these quantities are treated as random variables and estimated simultaneously, and their corresponding uncertainty is included in the resulting inferences.
-In this paper, we develop a computationally-efficient Gibbs sampler for estimation and apply the model to real and simulated data.
-The rest of this work is organized as follows.
-We first review relevant literature related to rank-clustering in Section 2.
+To induce rank-clusters, we place a novel spike-and-slab fusion prior on the object-specific parameters of BTL distributions.
+In contrast to existing work related to rank-clustering in the literature, our model requires neither the parameter order nor the number or size of rank-clusters to be known in advance.
+Instead, these quantities are treated as random variables and estimated simultaneously so that their corresponding uncertainty is naturally reflected in the resulting inferences.
+The rest of the paper is organized as follows.
+We first review literature related to rank-clustering in Section 2.
 Then, we propose the Partition-based Spike-and-Slab Fusion prior and apply it to a BTL model for ordinal data in Section 3.
-We develop a computationally-efficient Gibbs sampler based on reversible jump Markov chain Monte Carlo in Section 4 and demonstrate its accuracy on simulated data.
-To demonstrate a wide variety of methodological values of our proposed framework, Section 5 applies the model to four real datasets: (i) complete rankings of sushi preferences provided by Japanese adults, (ii) partial rankings of 2021 Minneapolis mayoral candidates expressed by voters in a ranked choice election, (iii) complete and partial rankings of policy options from Eurobarometer 34.1, a survey which measures various European attitudes, and (iv) pairwise basketball game outcomes from the 2020-21 season of the National Basketball Association.
+We develop a computationally-efficient Gibbs sampler based on reversible jump Markov chain Monte Carlo and demonstrate its accuracy on simulated data in Section 4.
+To demonstrate a wide variety of methodological benefits of our proposed framework, in Section 5, we apply the model to four real datasets: (i) complete rankings of sushi preferences provided by Japanese adults, (ii) partial rankings of 2021 Minneapolis mayoral candidates expressed by voters in a ranked choice election, (iii) complete and partial rankings of policy options from Eurobarometer 34.1, a survey which measures various European attitudes, and (iv) pairwise basketball game outcomes from the 2023-24 season of the National Basketball Association.
 We conclude with a brief discussion in Section 6.
 2 Background
-Before considering existing literature related to this paper, it is helpful to review some basic terminology and notation for ranking data.
-Rankings are a type of ordinal preference data that denote a relative ordering of objects from best to worst (potentially allowing ties).
+Before reviewing ordinal comparisons literature, it is helpful to introduce some basic terminology and notation.
+Rankings are a type of ordinal preferences that denote a relative ordering of objects from best to worst, potentially allowing ties.
 We use the operator $\prec$ to denote a strict ordering of two objects; e.g., $A\prec B$ states that object $A$ is strictly preferred to $B$.
-In a ranking, an object’s rank is the place it receives in the ranking.
+An object’s rank is the place it receives in the ranking.
 [^1] Although some authors have drawn a distinction between the terms “ranking” and “ordering,” in this paper we choose to use solely the former in accordance with its popular usage.
@@
 We call incomplete rankings involving two objects (candidates in the above example) a pairwise comparison, and incomplete rankings involving more than two objects a groupwise comparison.
-Rankings may be both partial and incomplete; e.g., a top-3 ranking of mayoral candidates only among the 5 candidates from a specific political party.
-In the rest of this section, we will briefly review methods for estimating rank-clusters based on the Bradley-Terry-Luce and Mallows families of ordinal data models in turn.
-For a more thorough review of these standard model families, see e.g., Marden (1996); Alvo and Yu (2014).
+Rankings may be both partial and incomplete; e.g., it could be a top-3 ranking of mayoral candidates from a specific political party.
+Next, we briefly review methods for estimating rank-clusters based on the Bradley-Terry-Luce and Mallows families of ordinal data models in turn.
+For a more thorough review of these standard model families, see Marden (1996) and Alvo and Yu (2014).
 2.1 Methods based on Bradley-Terry-Luce Distributions
@@
 As a result, Jeon and Choi (2018) proposed a modified regularization penalty that may be applied to partial or complete rankings under the Plackett-Luce model.
-Parameter fusion methods for rank-clustering exhibit five distinct disadvantages: First, maximum likelihood estimation of models in the BTL family, even in their simplest forms, often suffers from numerical instability and slow computational speed.
+Relatedly, Hermes et al.
+(2024) consider sparse estimation of a Plackett-Luce model with object-level covariates under judge heterogeneity.
+In their setting, the number of heterogeneous preference groups and the group membership of each judge are assumed fixed and known.
+To improve efficiency of estimation across groups and predictive performance, they impose a lasso penalty on group-specific covariate coefficients and a simultaneous fused lasso penalty between each pair of group-specific covariate coefficients.
+We note that the setting studied by Hermes et al.
+(2024) is fundamentally different to ours, in that they assume (known) preference heterogeneity among the judges and the presence of object-specific covariates.
+Parameter fusion methods for rank-clustering exhibit four distinct disadvantages: First, maximum likelihood estimation of models in the BTL family, even in their simplest forms, often suffers from numerical instability and slow computational speed.
 As a result, numerous authors have proposed complex algorithms to improve estimation accuracy or speed (Hunter et al., 2004; Maystre and Grossglauser, 2015; Turner et al., 2020; Nguyen and Zhang, 2023).
@@
 Fourth, prior knowledge on the amount and size of rank-clusters cannot be directly incorporated into the frequentist framework: Although the penalty parameter influences estimation of rank-clusters, the specific meaning of various possible choices is not directly interpretable.
-Fifth, to our knowledge only Jeon and Choi (2018) have formulated a parameter fusion-based method for ordinal data types beyond simple pairwise comparisons.
-As a result, only their parameter fusion method is available to estimate rank-clusters from complete or partial ranking data.
-Many of these disadvantages may be addressed using Bayesian methods.
-For example, Bayesian counterparts of the lasso and fused lasso—the Bayesian lasso (Park and Casella, 2008) and Bayesian fused lasso (Casella et al., 2010)—allow for valid uncertainty quantification and incorporation of prior knowledge via the selection of interpretable hyperparameters, as opposed to opaque penalty parameters.
-Bayesian methods for shrinkage and variable selection can be broadly categorized into two classes, continuous shrinkage and spike-and-slab priors.
-Examples of continuous shrinkage priors include the Bayesian Lasso (i.e., Laplace; Park and Casella (2008)), Bayesian Fused Lasso (Casella et al., 2010), t-distribution (Song and Cheng, 2020), Normal-Exponential-Gamma (Griffin and Brown, 2005; Shimamura et al., 2019), Normal-Gamma (Griffin and Brown, 2010), horseshoe (Carvalho et al., 2010), and Dirichlet-Laplace (Bhattacharya et al., 2015).
-No continuous shrinkage priors place positive probability on coefficients (or their differences) being precisely zero, which would encourage sparsity.
-Thus, parameter fusion must be performed via thresholding the posterior distribution, which is often ad-hoc (Porwal and Rodriguez, 2021).
-This leads us to consider priors of the latter class.
-Spike-and-slab priors (Mitchell and Beauchamp, 1988; George and McCulloch, 1997; Ishwaran and Rao, 2005) assign weight to both a point-mass at 0 (“spike”) and a continuous density function (“slab”).
-Although the specific formulations of these priors vary, they allow us to estimate which parameters are precisely zero in a probabilistic framework.
-Thus, spike-and-slab priors can be used for parameter fusion and uncertainty quantification.
-However, we are aware of only one existing modification of this prior class for parameter fusion: Wu et al.
-(2021) apply spike-and-slab to the differences in successive parameters in the linear regression setting.
+Many of these disadvantages may be addressed using spike-and-slab priors, a Bayesian approach to variable selection (Mitchell and Beauchamp, 1988; George and McCulloch, 1997; Ishwaran and Rao, 2005).
+Spike-and-slab priors assign weight to both a point-mass at 0 (“spike”) and a continuous density function (“slab”).
+Although the specific formulations of these priors vary, they estimate parameters which are precisely zero in a probabilistic framework that incorporates prior knowledge via interpretable hyperparameters, as opposed to opaque penalty parameters.
+However, we are aware of only one variant of this prior class for parameter fusion: Wu et al.
+(2021) apply spike-and-slab to differences in successive parameters in a linear regression.
 In their method, the order of parameters from least to greatest in coefficient value must be known in advance (as in the fused lasso).
-This is not practical in the ordinal data setting because the parameter order is equivalent to the overall ranking, whose estimation is a primary goal in the canonical preference data problem.
+This is not practical in the canonical ordinal data setting because the parameter order is equivalent to the overall ranking, whose estimation is a primary goal.
 Thus, no Bayesian parameter fusions methods exist which may be directly applied to ordinal data analyses with rank-clustering.
+Alternatively, one may consider the class of continuous shrinkage priors, which include Bayesian variants of the lasso (Park and Casella, 2008) and fused lasso (Casella et al., 2010) among others (e.g., Griffin and Brown (2005); Carvalho et al.
+(2010); Bhattacharya et al.
+(2015)).
+However, continuous shrinkage priors do not place positive probability on coefficients (or their differences) being precisely zero.
+Thus, parameter fusion must be performed via thresholding the posterior distribution, which is often ad-hoc (Porwal and Rodriguez, 2021) and will not be considered in this work.
 2.2 Methods based on Mallows Distributions
@@
 As will be shown in Section 3, our proposed model avoids all three issues by incorporating parameter fusion in the continuously-parameterized Bradley-Terry-Luce model family.
-3 Methods
+3 The Rank-Clustered Bradley-Terry-Luce Model
 In this section, we first develop a novel spike-and-slab prior for parameter fusion based on partitions.
@@
 $K=1$$\{4^{0},4^{0},4^{0},4^{0},4^{0},4^{0},4^{0},4^{0}\}$
-[…]
-[…]
+$K=2$$\{4^{0},4^{0},4^{0},4^{0},4^{1},4^{1},4^{1},4^{1}\}$
+$K=4$$\{4^{0},4^{0},4^{1},4^{1},4^{2},4^{2},4^{3},4^{3}\}$
 $K=8$$\{4^{0},4^{1},4^{2},4^{3},4^{4},4^{5},4^{6},4^{7}\}$
@@
 5 Applications
-In this section, we apply our model to four real datasets involving ordinal comparisons.
-These four applications were chosen to highlight the applicability of our method to various ordinal data types and domain areas, which are summarized in Table 2.
-The data sets are comprised of sushi preferences of Japanese adults (Kamishima, 2003), policy preferences of respondents from Great Britain in a Eurobarometer survey (Reif and Melich, 1993), ranked-choice votes in a Minneapolis mayoral election (Minneapolis Elections and Voter Services, 2021), and pairwise game outcomes among teams in the National Basketball Association (National Basketball Association, 2021).
-Section Setting Methodological Value
-5.1 Sushi preferences in Tohoku Rank-clustering with complete rankings; comparing inferred overall ranking with that of the Clustered Mallows Model.
-5.2 Minneapolis mayoral election votes Rank-clustering with top-3 partial rankings among 17 candidates; interpretable overall ranking captures the winner’s mandate in ranked-choice elections; comparing inferred overall ranking with those from a standard BTL model and two election procedures.
-5.3 Eurobarometer survey policy preferences Rank-clustering with partial rankings; inferred overall ranking permits indistinguishability among options to aid policymakers.
-5.4 Basketball game outcomes Rank-clustering with pairwise comparison data; inferred overall order of captures similarly-performing teams.
-Table 2: Summary of applications.
+In this section, we apply the Rank-Clustered BTL model to four real datasets involving ordinal comparisons.
+These four applications were chosen to highlight the applicability of our method to various ordinal data types and domain areas and illustrate methodological values of our approach which are summarized in Table 2.
+The data sets are comprised of sushi preferences of Japanese adults (Kamishima, 2003), policy preferences of respondents from Great Britain in a Eurobarometer survey (Reif and Melich, 1993), ranked-choice votes in a Minneapolis mayoral election (Minneapolis Elections and Voter Services, 2021), and pairwise game outcomes among teams in the United States National Basketball Association (National Basketball Association, 2024).
+Setting Data Type Methodological Value
+5.1 Sushi preferences in Tohoku Complete rankings of 10 sushi types Rank-clusters sushi types by preferences. Comparing inferred overall ranking with that of the Clustered Mallows Model.
+5.2 Minneapolis mayoral election votes Top-3 partial rankings of 17 candidates Interpretable overall ranking captures the winner’s mandate in ranked-choice elections.
+Comparing inferred overall ranking with those from a standard BTL model and two election procedures.
+5.3 Eurobarometer survey policy preferences Partial rankings of 7 policy options Inferred overall ranking permits identification of similarly preferred options to aid policymakers.
+5.4 Basketball game outcomes Pairwise comparisons (game winners) among 30 teams Inferred overall order captures similarly-performing teams. Setting with limited information and low signal.
+Table 2: Summary of applications by subsection.
 5.1 Sushi Preferences in Tohoku
-We first study complete preference rankings of 10 common sushi types based on a benchmarking dataset by Kamishima (2003).
-To allow our results to be comparable with an analysis of the sushi data in the related work of Piancastelli and Friel (2024), we analyze the preferences of survey respondents who lived in Japan’s Tohoku region until at least 15 years of age.
-In total, there were 280 respondents providing complete rankings of 10 sushi types.
+We first study complete preference rankings of 10 sushi types from a benchmarking dataset by Kamishima (2003).
+To allow our results to be comparable with an analysis of the sushi data by Piancastelli and Friel (2024), we analyze the preferences of survey respondents who lived in Japan’s Tohoku region until at least 15 years of age.
+There were 280 such respondents.
 We fit a Rank-Clustered BTL distribution to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering but permit a wide variety of outcomes.
 Figure 5 displays posterior rank-clustering probabilities (left) and parameter posteriors (right).
-Additional results and convergence diagnostics provided in the supplementary materials.
-Figure 5: Primary results of an analysis of preference rankings of 10 common sushi types in a survey of 280 Japanese adults.
-Posterior rank-clustering probabilities (left) and posterior distributions of sushi-specific worth parameters (right).
-Sushi types are ordered according to their posterior median worth.
-Based on the left panel in Figure 5, fatty tuna appears to be the strictly most preferred sushi type based on this population, followed by tuna and shrimp rank-clustered with high posterior probability in second place.
+Additional results and convergence diagnostics are provided in the supplementary materials.
+Figure 5: Primary results from Rank-Clustered BTL analysis of Tohoku sushi data.
+Left: Posterior rank-clustering probabilities.
+Main diagonal displays posterior median estimate of worth parameter after normalization.
+Red squares indicate maximum a posteriori rank-clusters.
+Right: Posterior distributions of sushi-specific worth parameters.
+Sushi types are ordered according to posterior median worth.
+Based on the left panel in Figure 5, fatty tuna appears to be strictly most preferred in this population, followed by tuna and shrimp rank-clustered in second place.
 Salmon roe and sea eel exhibit high posterior probability of rank-clustering, as do sea urchin, tuna roll, and squid; these two groups may themselves be rank-clustered.
@@
 Conversely, if the winner is a singleton first-place rank-cluster—clearly ranked above all other candidates—their mandate may be considered strong.
-We fit a Rank-Clustered BTL to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage few of rank-clusters.
+We fit a Rank-Clustered BTL to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage few rank-clusters.
 Figure 6 displays posterior rank-clustering probabilities (left) and parameter posteriors (right).
-Additional results and convergence diagnostics provided in the supplementary materials.
-Figure 6: Primary results of an analysis of 1000 valid votes in the 2021 Minneapolis mayoral election.
-Posterior rank-clustering probabilities (left) and posterior distributions of candidate-specific worth parameters (right).
+Additional results and convergence diagnostics are provided in the supplementary materials.
+Figure 6: Primary results from Rank-Clustered BTL analysis of mayoral votes.
+Left: Posterior rank-clustering probabilities.
+Main diagonal displays posterior median estimate of worth parameter after normalization.
+Red squares indicate maximum a posteriori rank-clusters.
+Right: Posterior distributions of candidate-specific worth parameters.
 In Figure 6, candidates are ordered by their posterior median estimate of worth.
@@
 We note that Frey is not rank-clustered with other candidates with high posterior probability, suggesting a relatively strong mandate.
-Cluster 2 consists of Kate Knuth and Sheila Nezhad, both non-incumbent DFL candidates with substantial political backing and fundraising.
-Clusters 3 and 4 consist of AJ Awed and Clint Conner, respectively, who are each DFL candidates with less experience and political backing than candidates in previous rank-clusters.
-Cluster 5 consists of Laverne Turner, a GOP candidate, and Cluster 6 consists of Carney Jr., Harcus, Globus, and Nelson.
+Cluster 2 consists of Kate Knuth and Sheila Nezhad, both female, non-incumbent DFL candidates.
 Last, Cluster 7 consists of 6 candidates with minimal support.
-Figure 7 compares point estimates of rank for each candidate among four methods.
+Figure 7 compares point estimates of rank for each candidate across four methods.
 Figure 7: Comparison of estimated rank for each candidate across four aggregation methods: Ranked Choice, First-Past-the-Post (FPP), BTL, and Rank-Clustered BTL (RC BTL).
@@
 The first and second rows display assigned ranks from ranked choice and “first-past-the-post” (FPP) election procedures, respectively.
-[^3] Since ranked choice and FPP elections use deterministic procedures, we use the term “assigned rank”.
 We calculate FPP ranks by ordering candidates by the number of first place votes he/she received (ignoring all second and third place votes).
 [^3] If the actual election had utilized FPP tabulation, results may have been different based on the differing voter strategies encouraged by ranked choice and FPP elections.
-The third and fourth rows display estimated ranks from a standard Bayesian BTL and our Rank-Clustered BTL, respectively.
-Estimated ranks are based on maximum a posteriori estimates.
+The third and fourth rows display maximum a posteriori ranks from a standard Bayesian BTL and our Rank-Clustered BTL, respectively.
 Frey wins the election in all methods.
 The BTL and Rank-Clustered BTL models roughly reflect the deterministic algorithms, although we notice some swaps in candidate ranks which may be attributed to differences between first place and second or third place votes.
-For example, Conner received fewer first place votes than Turner, but far more second and third place votes.
+For example, Conner received fewer first place votes than Turner, but far more second and third place votes (see supplementary materials for vote totals).
 As a result, deterministic algorithms rank Turner above Conner, while the BTL model takes into account the additional preference information and ranks Conner above Turner.
@@
 We seek to identify a population-level ordering of the priorities that accounts for potential equality or indistinguishability among the options based on the survey data.
-We note that these data were previously studied by Wang et al.
-(2017) in a mixed-membership model.
-Our analysis, although a simplification of the diverse population’s heterogeneous preferences, provides a simpler interpretation to policy-makers interested in understanding holistic attitudes of their constituent population.
+These data were previously studied by Wang et al.
+(2017) with a mixed-membership model to learn about heterogeneity of opinions among survey respondents.
+Our analysis, although a simplification of the diverse population’s heterogeneous preferences, provides a simpler interpretation to policy-makers interested in understanding rank-ordering of policy preferences.
 We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=2$ to encourage rank-clustering.
@@
 Additional results and convergence diagnostics provided in the supplementary materials.
-Figure 8: Primary results of an analysis of respondents from Great Britain to question 28 in the Eurobarometer 34.1 survey.
-Posterior rank-clustering probabilities (left) and posterior distributions of survey option-specific parameters (right).
-We observe that policy option 2 (information campaigns) is strictly preferred to the rest among the population of survey respondents from Great Britain, whereas options 5 (funding research) and 3 (legal penalty) are rank-clustered last.
+Figure 8: Primary results from Rank-Clustered BTL analysis of Eurobarometer 34.1 data.
+Left: Posterior rank-clustering probabilities.
+Main diagonal displays posterior median estimate of worth parameter after normalization.
+Red squares indicate maximum a posteriori rank-clusters.
+Right: Posterior distributions of policy-specific worth parameters.
+Policy option 2 (information campaigns) is strictly preferred to the rest among the population of survey respondents from Great Britain, whereas options 5 (funding research) and 3 (legal penalty) are rank-clustered last.
 The results indicates to policymakers that respondents in Great Britain strongly prioritize Option 2 in comparison to the rest, while pairs of Options 1 and 4 and Options 3 and 5, are, respectively, indistinguishable within each pair, with 1 and 4 being strongly preferred to 3 and 5.
-By rank-clustering similarly-preferred options, interpretation of the population’s preferences is simplified for policymakers.
-5.4 2020-21 National Basketball Association Game Outcomes
-Finally, we analyze outcomes of 1,249 pairwise games in the 2020-21 season of the National Basketball Association (NBA).
-In this season, there were 30 teams, who each played every other team at least twice; 3 pairs of teams played 11 matches.
+By rank-clustering similarly-preferred options, interpretation of constituent preferences is simplified for policymakers.
+5.4 2023-24 National Basketball Association Game Outcomes
+Last, we analyze outcomes of 1,230 games from the 2023-24 season of the National Basketball Association (NBA) of the United States of America (National Basketball Association, 2024).
+In this season, 30 teams each played 82 games, including between 2 and 5 games against every other team.
 We seek to estimate an overall ranking of teams that allows for potential equality in ranking.
-We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=0.1$ to encourage rank-clustering given the limited ordinal comparison data provided by pairwise matchups.
+We fit a Rank-Clustered Bradley-Terry-Luce model to the data with $a_{\gamma}=5$, $b_{\gamma}=3$, and $\lambda=1$ to encourage rank-clustering given the limited ordinal comparison data provided by pairwise matchups.
 Figure 9 displays posterior rank-clustering probabilities (left) and parameter posteriors (right).
-Additional results and convergence diagnostics provided in the supplementary materials.
-Figure 9: Posterior rank-clustering probabilities (left) and posterior distributions of basketball teams (right) based on the results of all 2020-21 regular season games.
+Additional results and convergence diagnostics are provided in the supplementary materials.
+Figure 9: Primary results from Rank-Clustered BTL analysis of 2023-24 NBA data.
+Left: Posterior rank-clustering probabilities.
+Main diagonal displays posterior median estimate of worth parameter after normalization.
+Right: Posterior distributions of team-specific worth parameters.
 In this setting, the Rank-Clustered BTL model estimates an ordering of professional basketball teams with uncertain rank-clustering patterns.
-Note than in the left panel of Figure 9, the squares on the main diagonal are black (indicating probability 1) by default since each team is inherently rank-clustered with itself.
-An 80% posterior credible interval indicates that there are between 4 and 6 rank-clusters; every team has less than 0.015 probability of belonging to a singleton rank-cluster.
-However, rank-clusters in this setting are less certain than in previous examples.
-Most clearly, 5 teams (Detroit Pistons, Oklahoma City Thunder, Cleveland Cavaliers, Orlando Magic, and the Houston Rockets) appear to be rank-clustered for last place.
-Among these 5 teams, pairwise rank-clustering probabilities are at least 0.75.
-Instead of ranking these teams with granularity, we may simply infer by our model that they were each approximately the worst teams of the league in this season.
-Conversely at the top of the overall ranking, we observe that the top 5 teams (Utah Jazz, Phoenix Suns, Philadelphia 76ers, Milwaukee Bucks, and the Brooklyn Nets) all have a high posterior probability of rank-clustering.
-Although the Milwaukee Bucks were the ultimately this season’s league champions, our model suggests that these 5 teams were of roughly indistinguishable ability based on the game data.
-These summaries from our model, despite not accounting for the complexities of the sport, provide useful and interpretable summaries of the teams’ abilities across the season.
+Uncertain rank-clustering may result from two aspects of this application.
+First, pairwise comparisons provide little information in relation to partial or complete rankings, by construction.
+Second, game outcomes provide low signal measurements of team ability (Baumer et al., 2023).
+That is because many factors influence game outcomes, such as skill, home advantage, injuries, roster changes, and luck (Cai et al., 2019).
+Consistent with the low signal and limited information setting, an 80% posterior credible interval indicates that there are between 6 and 9 rank-clusters.
+Every team has less than 0.037 posterior probability of belonging to a singleton rank-cluster.
+As seen in the left panel of Figure 9, 4 teams (Boston Celtics, Oklahoma City Thunder, Denver Nuggets, and Minnesota Timberwolves) appear to be rank-clustered for first place.
+Based on regular season data alone, our model suggests that these 4 teams were of roughly indistinguishable ability.
+Conversely, we observe that 6 teams (Toronto Raptors, San Antonio Spurs, Portland Trail Blazers, Charlotte Hornets, Washington Wizards, and Detroit Pistons) all have a high posterior probability of rank-clustering in last place.
+Instead of reporting the uncertain ranking of these teams with some granularity, we recommend to infer that these teams were the worst teams of the league in this season.
+These rank-clusters, despite not accounting for the complexities of the sport, provide useful and interpretable summaries of the teams’ abilities across the regular season.
+A similar analysis could be used in the future to predict postseason performance.
 6 Discussion
 In this paper, we proposed the Rank-Clustered Bradley-Terry-Luce model for estimating an overall ranking of objects with rank-clusters.
-The model employs the Bradley-Terry-Luce (BTL) family of distributions for ordinal comparison data and estimates model parameters in a Bayesian framework under the proposed Partition-based Spike-and-Slab (PSSF) prior.
-The model requires no pre-specification of the number or size of the rank-clusters (as in Piancastelli and Friel (2024)), nor specification of lasso-based penalty parameters (as in Masarotto and Varin (2012); Jeon and Choi (2018)).
-In a simulation study, we demonstrated the model’s ability to accurately and consistently estimate the relative worth of objects in a collection while simultaneously estimating rank-clusterings over objects.
+The model employs the Bradley-Terry-Luce (BTL) family of distributions for ordinal comparisons.
+We proposed Partition-based Spike-and-Slab Fusion (PSSF) prior to estimates model parameters in a Bayesian framework.
+The model requires neither pre-specification of the number or size of the rank-clusters (improving upon Piancastelli and Friel (2024)), nor specification of lasso-based penalty parameters (improving upon Masarotto and Varin (2012); Jeon and Choi (2018); Hermes et al.
+(2024)).
+In a simulation study, we demonstrated the model’s ability to accurately and consistently estimate the relative worth of objects in a collection while simultaneously estimating rank-clusters.
 We used Rank-Clustered BTL on four real datasets under different types of ordinal comparison data.
@@
 Estimation of parameters under this model requires reversible jump MCMC.
-To overcome potentially slow or computationally-burdensome estimation in this setting, we proposed a computationally efficient Gibbs sampler that alternates between updating the partition of objects (based on the seminal work of Green (1995) and updating object-level worth parameters by adapting a data augmentation trick for standard Plackett-Luce models by Caron and Doucet (2012) and later adapted for Plackett-Luce mixtures by Mollica and Tardella (2017).
-A useful benefit of estimating parameter values and clusters in a single Bayesian framework is the avoidance of issues associated with selective inference (Taylor and Tibshirani, 2015) or more colloquially, double dipping (Kriegeskorte et al., 2009).
+To overcome potentially slow or computationally-burdensome estimation in this setting, we proposed a computationally efficient Gibbs sampler.
+The sampler alternates between updating the partition of objects, based on the seminal work of Green (1995), and updating object-level worth parameters following a data augmentation trick for standard Plackett-Luce models by Caron and Doucet (2012) that was later adapted for Plackett-Luce mixtures by Mollica and Tardella (2017).
+The proposed PSSF prior requires selecting hyperpriors for partitions, $f_{G}$, and the continuous values for each unique parameter, $f_{\nu}$.
+In this work, we specified $f_{G}\propto\text{Poisson}(K_{g}|\lambda)$ to be intentionally vague over the large space of partitions and $f_{\nu}\propto\text{Gamma}(a_{\gamma},b_{\gamma})$ based on conjugacy.
+However, alternative hyperpriors are available.
+A Negative Binomial or Beta Negative Binomial distribution for $f_{G}$ may be more appropriate when stronger prior knowledge of $K_{g}$ is available.
+If PSSF were to be applied to linear regression for parameter fusion, a Normal or $t-$distribution may be substituted for $f_{\nu}$.
+A useful benefit of estimating parameter values and clusters in a single Bayesian framework is the avoidance of issues associated with selective inference (Taylor and Tibshirani, 2015) or, more colloquially, double dipping (Kriegeskorte et al., 2009).
 Selective inference occurs when the same data is used twice in the process of model selection and/or estimation, e.g., to estimate some latent structure underlying the data and subsequently to estimate parameters conditional on that estimated structure.
@@
 As such, our parameter estimates incorporate uncertainty across the posterior distributions of both the rank-clustering structure and the specific parameter values.
-Results from Rank-Clustered BTL models are useful in a variety of contexts.
+Results from Rank-Clustered BTL models are useful in a variety of inferential contexts.
 As noted in other fusion literatures on rankings, estimated overall rankings may be easier to understand and interpret when rank-clusters of objects are identified, as rank-clusters lead to fewer rank levels of objects to distinguish (Masarotto and Varin, 2012).
@@
 In the former, the PSSF prior could be applied to preference learning via BTL distributions that incorporate covariates (e.g., Baldassarre et al.
-(2023); Gormley and Murphy (2010); Chapman and Staelin (1982)).
+(2023); Gormley and Murphy (2010); Chapman and Staelin (1982); Hermes et al.
+(2024)).
 In that case, the prior may be modified to permit covariate parameter estimation in addition to rank-clustering.
@@
 (2019)).
+Third, we notice that the PSSF prior bears some resemblance to a Dirichlet process prior (Escobar and West, 1995).
+The connection between Bayesian nonparametrics and Bayesian parameter fusion requires further study.
 The proposed Rank-Clustered BTL model accurately estimates rank-clusters, permitting complex summaries beyond the traditional overall ranking and allowing for improved interpretability of the results.
@@
 Appendix A.4 Additional Results from Section 5.4
-Figure 19 displays stacked bar charts of the season record of each NBA team across the 2020-21 season.
-Figure 19: Stacked bar charts of ranks received by each NBA team across the 2020-21 season. Winning = rank 1; losing = rank 2.
+Figure 19 displays stacked bar charts of the season record of each NBA team across the 2023-24 season.
+Figure 19: Stacked bar charts of ranks received by each NBA team across the 2023-24 season. Winning = rank 1; losing = rank 2.
 Figures 20 and 21 contain trace plots for $K$ and $\omega$ after burn-in for each chain. We find the trace plots to demonstrate satisfactory mixing and convergence.
-Figure 20: Trace plot of K in the 2020-21 NBA season analysis
-Figure 21: Trace plot of $\omega$ in the 2020-21 NBA season analysis
+Figure 20: Trace plot of K in the 2023-24 NBA season analysis
+Figure 21: Trace plot of $\omega$ in the 2023-24 NBA season analysis